# King’s College London, U.K

### 13 Mar 2024 v2

@@ Abstract
We demonstrate that Assembly Theory, pathway complexity, the assembly index, and the assembly number are subsumed and constitute a weak version of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity reliant on an approximation method based upon statistical compression, their results obtained due to the use of methods strictly equivalent to the LZ family of compression algorithms used in compressing algorithms such as ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT’s assembly index and their use had already been reported in successful application to separating organic from non-organic molecules, and the study of selection and evolution. Here we exhibit and prove the connections and full equivalence of Assembly Theory to Shannon Entropy and statistical compression, and AT’s disconnection as a statistical approach from causality. We demonstrate that formulating a traditional statistically compressed description of molecules, or the theory underlying it, does not imply an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by selection and evolution, when lacking in logical consistency and empirical evidence. We argue that in their basic arguments, the authors of AT conflate how objects may assemble with causal directionality, and conclude that Assembly Theory does not explain selection or evolution beyond known and previously established connections, some of which are reviewed here.
@@
assembly theory, assembly index, pathway complexity, assembly number, algorithmic complexity, algorithmic probability, BDM, lossless compression, LZ compression, LZ77, LZ78, LZW, selection, evolution.
@@ 2.1 The assembly Index is a compression algorithm of the LZ family
In contrast, more robust approximations to $\mathbf{K}$, capable of capturing blocks and other causal content within an object, have been proposed for purposes ranging from exploring cause-and-effect chains to quantifying object memory and characterising process content [16]. These advanced measures have found application in the same contexts and domains as Assembly Theory, encompassing tasks like distinguishing organic from non-organic molecules [17], investigating potential connections to selection and evolution [6, 7], the detection of bio and techno signatures [18, 19], and explorations into causality [20, 21]. Importantly, when applied to the data employed as evidence by AT, the more sophisticated measures consistently outperform the assembly index (see Suplementary Information and [12]). Thus, it becomes evident that Assembly Theory and its assembly index represent a considerably constrained version of compression algorithms, and a loose upper bound of $\mathbf{K}$.
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The emphasis on the purported requirement of a Turing machine in the context of algorithmic complexity is false and a misdirected concern.
@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in AT [31]. This Venn diagram shows how Assembly Theory (AT) is connected and subsumed within algorithmic complexity or $\mathbf{K}$ and even within statistical compression as proven in this paper (see Sup. Inf.). B: Causal transition graph from a Turing machine with number 3019 (in Wolfram’s enumeration scheme [32]) with an empty initial condition found by using a computable method (e.g. CTM, [33]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process, as the authors of AT have claimed. Turing’s motivation was to explore what could be automated in the causal mechanisation of operations, for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram.
@@
Nevertheless, the question of random versus non-random selection and evolution in the context of approximations to algorithmic complexity, including copy counting, was experimentally tested, empirically supported, and reported before in [6] following proper scientific practice, including a basic literature search and review, control experiments (comparison to other measures), and validation against existing knowledge in genetics and cell biology.
These circular arguments, lacking a foundation in chemistry, biology, or empirical evidence, lead the authors to propose the concept of “assembly time”. They suggest a time-scale separation between molecule production (assembly time) and discovery time, aiming to unify physics and biology. Time was already fundamental in evolutionary theory. Assembly Theory revisits existing complexity science theories rather than offering any meaningful new insight (see Fig. 2).
Figure 2: A timeline of results in complexity science relevant to the claims of Assembly Theory. Assembly Theory borrows almost verbatim ideas from previous decades, renaming LZ77/LZ78 an ‘assembly index’, dictionary trees ‘assembly pathways’ and taking algorithmic probability as their evolution equation, overlooking 50-80 years of developments and results. Assembly Theory is identical to Logical Depth in its definition of the characterisation of complexity by number of steps, the only difference being that the authors’ claim it is not based on compression and therefore $\mathbf{K}$, as Logical Depth is. The fact of the matter, however, is that it is. The assembly index is equivalent to LZ lossless compression algorithms.
@@ 3 Conclusions
We have shown how the assembly index results can be explained in the context of a dictionary-based compression algorithm that belongs to the LZ family. The most popular examples used by the authors of Assembly Theory, such as ABRACADABRA and BANANA are widely used to teach LZ77/LZ78 compression at under-grad level. The results obtained by Assembly Theory to date have been reported in previous literature in the context of algorithmic information theory [17, 6]. The concepts, results, index, and threshold, as reported in [1, 35, 4, 2], follow directly from earlier work on algorithmic complexity.
As shown before, on the one hand, indexes, including other popular lossless statistical compression algorithms, produce similar or better results on the same data and from multiple representations of the same chemical data [12] as the one used by Assembly Theory.
On the other hand, here we have proven that the main components and substance of Assembly Theory are based upon and subsumed into the theory of algorithmic complexity. Based on their high and low copy number feature driving pathway complexity and their assembly number, an object with a low assembly index will have low Entropy and high LZ compressibility, and an object with a high assembly index will have high Entropy and low LZ compressibility with respect to the size of the object in the same proportions and driven by the same copy-counting mechanism (on which every other popular statistical compression is also based upon). This means that Assembly Theory is a weak approximation to algorithmic complexity. The latter has more discriminatory power because it can tell apart cases of high assembly index with low copy number, high Entropy, and low LZ compressibility but with potentially causal compressible content [12] that can be described by short rewriting rules of the same type of but a non-trivial nature than those only taken into account by Assembly Theory. Thus, algorithmic complexity is more profoundly connected to causality [36, 37, 20, 38]. The claim that it can only be applied to strings is also false, as it has been applied before to images [39], networks [7, 40, 41, 42], vectors, matrices, and tensors in multiple dimensions [43, 44, 45]; and even chemical structures [17].
@@
We agree with the authors on the importance of understanding the origin and mechanisms of the emergence of an open-ended generation of novelty [47, 38]. However, these efforts by the authors are undermined by hyperbolic claims through their university press releases, journals, online media, and the press.
The suggestions made by Assembly Theory that people may resonate with, only highlight the working hypothesis that information theory, algorithmic complexity, and algorithmic probability underpin the processes and building blocks of physical and living systems on which all these methods and frameworks have been heavily inspired upon [48, 49], playing an important role in chemical processes [17], nature and biology as suggested before [17, 50, 6, 39].
@@ A.3.1 Assembly index defines an encoding scheme
    The first equation is by construction of $G$. The second is a direct consequence of the first and [54, Theorem $1$]. ∎
@@ A.3.3 Assembly index is a version of LZ compression
    In the same manner, Theorem 9 demonstrates that “simpler” minimum rooted assembly subspaces lead to more compressibility. For example, the reader is invited to note that due to the fact that $\gamma_{y}^{\max}=\gamma_{y}^{\min}$ would hold, a linear minimum rooted assembly subspace ${\Gamma^{*}}_{y}$, such as in [1, Figure 1(a)] or [51, Figure 8], results in a much tighter upper bound in Equation (10). This is because the exponential dependence on the assembly index to encode all possible DAG-like structures would drop: that is, $1\leq t\leq\mathbf{O}\left(\left|\gamma_{y}^{\max}\right|^{2}\,\left|B_{\Gamma}\right|\right)$ will hold instead of Equation (15), which in turn will be subsumed in the rightmost term of Equation (10) where one already has a quadratic dependence on $\left|\gamma_{y}^{\max}\right|$ or $c_{\Gamma}\left(y\right)$.

### 14 Mar 2024 v3

@@ Abstract
We demonstrate that Assembly Theory, pathway complexity, the assembly index, and the assembly number are subsumed and constitute a weak version of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity reliant on an approximation method based upon statistical compression, their results obtained due to the use of methods strictly equivalent to the LZ family of compression algorithms used in compressing algorithms such as ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT that were reported before in successful application to separating organic from non-organic molecules and in the context of the study of selection and evolution. We prove the connections and full equivalence of Assembly Theory to Shannon Entropy and statistical compression, and AT’s disconnection as a statistical approach from causality. We demonstrate that formulating a traditional statistically compressed description of molecules, or the theory underlying it, does not imply an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by selection and evolution, when lacking in logical consistency and empirical evidence. We argue that in their basic arguments, the authors of AT conflate how objects may assemble with causal directionality, and conclude that Assembly Theory does not explain selection or evolution beyond known and previously established connections, some of which are reviewed.
@@
assembly theory, assembly index, pathway complexity, assembly number, algorithmic complexity, algorithmic probability, CTM, BDM, lossless compression, LZ compression, LZ77, LZ78, LZW, Lempel–Ziv–Welch, logical depth, selection, evolution.
@@ 2.1 The assembly Index is a compression algorithm of the LZ family
In contrast, more robust approximations to $\mathbf{K}$, capable of capturing blocks and other causal content within an object, have been proposed for purposes ranging from exploring cause-and-effect chains to quantifying object memory and characterising process content [16, 17]. These more advanced measures have found application in the same contexts and domains explored by Assembly Theory, encompassing tasks like distinguishing organic from non-organic molecules [18], investigating potential connections to selection and evolution [6, 7], the detection of bio and techno signatures [19, 20], and explorations into causality [21, 22]. Importantly, when applied to the data employed as evidence by AT, the more sophisticated measures consistently outperform the assembly index (see Suplementary Information and [12]). Thus, it becomes evident that Assembly Theory and its assembly index represent a considerably constrained version of compression algorithms, and a loose upper bound of $\mathbf{K}$.
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The emphasis on the purported requirement of a Turing machine in the context of algorithmic complexity is a misdirected concern. A Turing machine is a synonym of ‘algorithm’ and an ‘algorithm’ is a synonym of a Turing machine, any rule or method that is algorithmic in nature, including all the computable methods in Assembly Theory, are algorithms, and, therefore, technically, Turing machines or programs that can run on a (universal) Turing machine. These facts are only a technicality unrelated to any putative advantage of Assembly Theory over other approaches to algorithmic complexity.
@@ 2.4 Lack of control experiments and absence of supporting empirical data beyond current domain knowledge
Turing’s motivation was to explore what could be automated in the causal mechanisation of operations (see Fig. 1), for example, stepwise by hand using paper and pencil. The shortest among all the computer programs of this type is an upper bound approximation to $\mathbf{K}$. In other words, $\mathbf{K}$ cannot be longer than the length of this diagram. The concept of pathway complexity and copy number instantiated by the assembly index, is a restricted version of statistical nature of a Turing machine or finite automaton.
Figure 1: A: The authors of AT have suggested that $\mathbf{K}$ would be proven to be contained in Assembly TheoryT [32]. This Venn diagram shows how Assembly Theory (AT) is connected and subsumed within algorithmic complexity or ($\mathbf{K}$) and within the group of statistical compression as proven in this paper (see Sup. Inf.). B: Causal transition graph of a Turing machine with number 3019 (in Wolfram’s enumeration scheme [33]) with an empty initial condition found by using a computable method (e.g. CTM, [34]) to explain how the block-patterned string 111000111000 was assembled step-by-step based on the principles of $\mathbf{K}$ describing the state, memory, and output of the process as a fully causal mechanistic explanation. A Turing machine is simply a procedural algorithm and any algorithm can be represented by a Turing machine. By definition, this is a mechanistic process, and as physical as anything else, not an ‘abstract’ or ‘unrealisable’ process.
@@
Nevertheless, the question of random versus non-random selection and evolution in the context of approximations to algorithmic complexity, including copy counting, was experimentally tested, empirically supported, and reported before in [6] following basic literature search and review, control experiments (comparison to other measures), and validation against existing knowledge in genetics and cell biology.
These circular arguments, lacking a foundation in chemistry, biology, or empirical evidence, lead the authors to propose the concept of “assembly time”. The suggestion is a time-scale separation between molecule production (assembly time) and discovery time, aiming to unify physics and biology. Time, however, was already fundamental in evolutionary theory. Assembly Theory revisits existing complexity science concepts as described in the timeline provided in Fig. 2.
Figure 2: A timeline of results in complexity science relevant to the claims and results from Assembly Theory. Assembly Theory renames several concepts such as dictionary trees ‘assembly pathways’, algorithmic probability and their reduction of combinatorial space. The assembly index and number are equivalent to LZ lossless compression algorithms with the full details provided in the Sup. Inf.
@@ 3 Conclusions
We have shown how the assembly index results can be explained in the context of a dictionary-based compression algorithm that belongs to the LZ family. The most popular examples used by the authors of Assembly Theory, such as ABRACADABRA and BANANA are traditionally used to teach LZ77/LZ78 compression in Computer Science courses at university level. Results obtained by Assembly Theory to date have been reported previously in the context of algorithmic information theory [18, 6]. The concepts, results, index, and threshold, as reported in [1, 36, 4, 2], follow directly from earlier work on algorithmic complexity and data compression.
Based on their high and low copy number feature driving pathway complexity and their assembly number, an object with a low assembly index will have low Entropy and high LZ compressibility, and an object with a high assembly index will have high Entropy and low LZ compressibility with respect to the size of the object in direct proportion and driven by the same copy-counting mechanism (on which every popular statistical compression is based upon).
In contrast, algorithmic (Kolmogorov) complexity and other computable approximations [17, 21] potentially have more discriminatory power because they may tell apart cases in which high assembly index values with low copy number, high Entropy, and low LZ compressibility are causally compressible by exploiting short rewriting describing the object beyond statistical identical copies.
Algorithmic complexity has been shown to be profoundly connected to causality [37, 38, 21, 39]. The claim that it can only be applied to strings is incorrect, as it has been applied before to images [40], networks [7, 41, 42, 43], vectors, matrices, and tensors in multiple dimensions [44, 45, 46]; and even chemical structures [18].
@@
We agree with the authors on the importance of understanding the origin and mechanisms of the emergence of an open-ended generation of novelty [48, 39]. However, these efforts by the authors are undermined by hyperbolic claims. The suggestions made by Assembly Theory that people may resonate with highlight the working hypothesis that information theory, algorithmic complexity, and algorithmic probability underpin the processes and building blocks of physical and living systems on which all these methods and frameworks have been heavily inspired upon [49, 50], playing an important role in chemical processes [18], nature and biology as suggested before [18, 51, 6, 40].
@@ A.3.1 Assembly index defines an encoding scheme
    The first equation is by construction of $G$. The second is a direct consequence of the first and [55, Theorem $1$]. ∎
@@ A.3.3 Assembly index is a version of LZ compression
    In the same manner, Theorem 9 demonstrates that “simpler” minimum rooted assembly subspaces lead to more compressibility. For example, the reader is invited to note that due to the fact that $\gamma_{y}^{\max}=\gamma_{y}^{\min}$ would hold, a linear minimum rooted assembly subspace ${\Gamma^{*}}_{y}$, such as in [1, Figure 1(a)] or [52, Figure 8], results in a much tighter upper bound in Equation (10). This is because the exponential dependence on the assembly index to encode all possible DAG-like structures would drop: that is, $1\leq t\leq\mathbf{O}\left(\left|\gamma_{y}^{\max}\right|^{2}\,\left|B_{\Gamma}\right|\right)$ will hold instead of Equation (15), which in turn will be subsumed in the rightmost term of Equation (10) where one already has a quadratic dependence on $\left|\gamma_{y}^{\max}\right|$ or $c_{\Gamma}\left(y\right)$.